OpenStudy (anonymous):

A falling skydiver is accelerating in the downward direction at 3.29 m/s/s. The mass of the skydiver (including parachute gear) is 67.2 kg. Determine the air resistance force on the skydiver (and accompanying parachute). I got 437.48N. Is that right?

OpenStudy (xishem):

\[\sum_{}^{}\vec F=m\vec a\]\[\vec F_D-\vec F_g=m\vec a\]\[\vec F_D=m\vec a + \vec F_g = (67.2kg)(-3.29\frac{m}{s^2})+(67.2kg)(9.81\frac{m}{s^2})=438N\]Yep, looks like you're right!

OpenStudy (anonymous):

AWESOME!! thanks so much